# GLS: Geometry-aware 3D Language Gaussian Splatting

### v1

@@ Abstract
Recently, 3D Gaussian Splatting (3DGS) has achieved significant performance on indoor surface reconstruction and open-vocabulary segmentation. This paper presents GLS, a unified framework of surface reconstruction and open-vocabulary segmentation based on 3DGS. GLS extends two fields by exploring the correlation between them. For indoor surface reconstruction, we introduce surface normal prior as a geometric cue to guide the rendered normal, and use the normal error to optimize the rendered depth. For open-vocabulary segmentation, we employ 2D CLIP features to guide instance features and utilize DEVA masks to enhance their view consistency. Extensive experiments demonstrate the effectiveness of jointly optimizing surface reconstruction and open-vocabulary segmentation, where GLS surpasses state-of-the-art approaches of each task on MuSHRoom, ScanNet++, and LERF-OVS datasets. Code will be available at https://github.com/JiaxiongQ/GLS.
@@ 1 Introduction
Surface reconstruction [18, 20, 54, 53, 8, 55] and open-vocabulary segmentation [34, 38, 56, 47, 19, 43] based on 3DGS [25], being widely applied in AR/VR [23] and embodied intelligence [6, 33] due to its capabilities of efficient training and real-time rendering. Recently, notable works have achieved significant progress in both areas.
@@
For surface reconstruction, SuGaR [18] proposes regularization terms to align Gaussians and scene surface, then uses Possion reconstruction [24] to extract mesh from Gaussians. For open-vocabulary segmentation, LangSplat [34] and OpenGaussian [43] successfully introduces SAM [27] and CLIP [22] to 3DGS. Gaussian grouping [47] utilizes a universal temporal propagation model DEVA [11] to obtain consistent object masks across views and propose a 3D regularization term to grouping Gaussians. However, these methods only focus on one task and suffer from unstable performance in complex indoor scenes as Fig. 1 shows.
@@
Figure 1: Indoor surface reconstruction and open-vocabulary segmentation. Shadow and high-light regions make state-of-the-art methods struggle in indoor scenes. Our proposed GLS jointly optimizes two tasks based on 3DGS and achieves much better results than OpenGaussian [43] and 2DGS [20].
@@
In this work, we aim to achieve efficient and robust indoor surface reconstruction and open-vocabulary segmentation based on 3DGS. Our goal has two main motivations. On the one hand, the 2D open-vocabulary supervision [27, 22] is naturally view-inconsistent, which easily results in noise body and blur boundary of the segmented object from Gaussians. The accurate scene surface consists of sharp and smooth object surfaces, illustrating that Gaussians are mainly distributed on objects and preserve the object boundaries. This property can make the object segmentation results from Gaussians cleaner and sharper. On the other hand, due to the complex materials and lighting conditions of the indoor scene, the shadow and high-light regions on texture-less and reflective objects always result in noisy surfaces. Fortunately, accurate object masks erase the interference details on the object’s surface. Hence, the object segmentation results can supply the smoothness prior to reducing the reconstruction noises of these objects. In general, the optimization goals of the two can be considered to be the same.
@@
Based on above motivations, we introduce GLS, which leverages complementary between surface reconstruction and open-vocabulary segmentation to boost the performance of both tasks. The framework is presented in Fig. 3. Specifically, we first introduce the normal prior [1] to regularize the surface normal estimated from rendered depth. Then we analyzed different situations of rendered depth under different normal errors and propose a regularization term to enhance the sharpness of the scene surface. To integrate the open-vocabulary information, we add Gaussian semantic features into vanilla 3DGS and then utilize the consistent object masks from DEVA and image features from CLIP to supervise them. In addition, we consider the clip features as the smoothness prior to strengthen the accuracy of texture-less and reflective surfaces. Finally, we adopt TSDF fusion [32] from rendered depth to extract scene mesh, then compute the similarity between input text embeddings and learned semantic embeddings to acquire object masks. We conduct extensive experiments on MuSHRoom [36], ScanNet++ [49] and LERF-OVS  [34, 26] datasets. As Fig. 2 shows, the superior performance of our model on both tasks demonstrates the effectiveness of connecting surface reconstruction and open-vocabulary segmentation in 3DGS.
In summary, our technical contributions can be listed as follows:
@@
    We design a novel framework based on 3DGS, by jointly optimizing surface reconstruction and open-vocabulary segmentation in complex indoor scenes.
@@
    We propose two regularization terms with the help of geometric and semantic cues, to facilitate the sharpness and smoothness of reconstructed scene surfaces and segmented objects.
@@
    Our method inherits the training and rendering efficiency of 3DGS and achieves state-of-the-art accuracy on surface reconstruction and open-vocabulary segmentation tasks.
Figure 2: Effect of joint optimization between indoor surface reconstruction and open-vocabulary segmentation. Our method significantly improves the quality of Gaussian semantic features and surface reconstruction.
Figure 3: Overview of our framework. We adopt geometric and semantic cues produced by generalizable models to jointly strengthen the reconstruction and segmentation quality of 3DGS. Our framework achieves accurate indoor surface reconstruction and open-vocabulary segmentation via effective regularization terms, which faithfully enable Gaussian primitives to distribute along the object surface.
@@ 2.1 Neural Rendering
Neural Radiance Fields (NeRF) [30] based on implicit representations have achieved remarkable advancements in novel view synthesis. Some NeRF-based approaches [2, 3, 4] concentrate on extending NeRF to more challenging scenes, while they are still limited by low training and rendering efficiency. Alternatively, other NeRF-based methods [39, 13, 35, 31, 50, 10] utilize explicit representations such as voxels, grids, point clouds and meshes. They faithfully reduce the large computational cost of neural networks. The recent technologies [25, 52, 28] of neural rendering based on alpha-blend rendering have further advanced the rendering speed and quality, by optimizing the attributes of the explicit 3D Gaussians. Our framework is built upon 3DGS and focuses on optimizing it to tackle both surface reconstruction and open-vocabulary segmentation tasks.
@@
Surface reconstruction from multi-view images [37, 15, 16, 17] is a challenging task in computer vision and graphics. Recently, extracting scene surfaces from Gaussian primitives has become a popular research topic. SuGaR [18] utilizes Poisson reconstruction [24] to extract mesh from sampled Gaussians and encourages Gaussians attach to the scene surface by several regularization terms, for enhancing reconstruction accuracy. However, due to the disorder of the Gaussians and the complexity of the scene, fitting Gaussians on geometric surfaces is challenging. 2DGS [20] replaces 3D Gaussian primitives with 2D Gaussian disks for efficient training. They also introduce two regularization terms of depth distortion and normal consistency to reconstruct a smooth scene surface. GOF [53] proposes the Gaussian opacity field, which extracts holistic mesh by identifying the scene level-set directly. PGSR [8] estimates the unbiased depth to disable the degeneration of the blending coefficient, and introduces the multi-view regularization term to achieve global-consistent reconstruction. DN-Splatter [40] and VCR-GauS [9] leverage normal or depth priors from generalizeable models [1, 5, 14], to learn effective surfel representations. GaussianRoom [44] integrates implicit SDF representation with explicit Gaussian representation, and utilizes normal and edge priors to improve the reconstruction accuracy.
Due to the lack of object attributes, these methods struggle to reconstruct sharp objects against complex lighting conditions of an indoor scene. To address these issues, we propose combining open-vocabulary segmentation and surface reconstruction. In contrast with previous related works [46, 41, 42], our method releases the need for ground-truth segmentation supervision and achieves highly efficient training speed.
@@ 4.1 Leveraging 2D Semantic and Geometric Cues
For open-vocabulary segmentation, we directly use the view-consistent segmentation results of DEVA [11] as the supervision of the rendered mask $M$. Specifically, given the rendered features $F$, we use an MLP layer to increase its feature dimension to the number of total categories first. The softmax function and standard cross-entropy loss $\mathcal{L}_{m}$ are adopted for final classification. To enhance the open-vocabulary capabilities of our model, we employ the CLIP features $\hat{F}$ to supervise $F$ by:
@@ 4.2 Semantic-feature Guided Normal Smoothing
Big surfaces like floors and desktops are generally over-smoothing in an indoor scene. Although $\mathcal{L}_{n}$ supply local smoothness for reconstruction, it still struggles in shadow and high-light regions of these surfaces as Fig. 5 shows. High-weight $\mathcal{L}_{n}$ causes an over-smoothing effect and then makes some small objects disappear, while low-weight $\mathcal{L}_{n}$ disables the local smoothness of rendered depth.
To resolve this issue, we propose introducing clip features to help smooth big surfaces, then we can employ $\mathcal{L}_{n}$ with low weight for protecting small objects. Essentially, clip features supply high-dimension object-level smoothness which can implicitly reduce the inference of big surfaces. Concretely, we first introduce SAM to select objects that occupy the top-$k$ area by the mask $M_{o}$. and design a novel regularization term $\mathcal{L}_{s}$ for smoothing them as:
@@ 4.4 Optimization
We set $\alpha_{n}=0.07$, $\alpha_{m}=0.3$, $\alpha_{clip}=1.0$, $\alpha_{d}=0.1$ and $\alpha_{s}=0.5$ in our experiments.
Table 1: Quantitative results of indoor surface reconstruction on MuSHRoom dataset. Our method outperforms other methods on most metrics. The best metrics are highlighted.
@@
DN-Splatter [40] Poisson 0.0402 0.0211 0.0306 0.8449 0.8590
Ours TSDF 0.0288 0.0254 0.0271 0.8640 0.8924
2DGS [20] $\times$ TSDF 0.1728 0.1619 0.1674 0.7113 0.3433
PGSR [8] $\times$ TSDF 0.4957 0.7672 0.6315 0.5566 0.1178
Ours $\times$ TSDF 0.0903 0.0703 0.0870 0.7830 0.5153
Table 2: Quantitative results of indoor surface reconstruction on ScanNet++ dataset. ’MC‘ means Marching Cube algorithm [29]. Our method achieves comparable performance and 30×30$\times$30 × training speed than MonoSDF [51]. The best metrics are highlighted.
@@
MonoSDF [51] MC 0.0536 0.0490 0.0513 0.9142 0.7124
DN-Splatter [40] Poisson 0.0977 0.0431 0.0704 0.8272 0.7094
Ours TSDF 0.0640 0.0272 0.0444 0.9064 0.8623
2DGS [20] $\times$ TSDF 0.2440 0.4362 0.3401 0.6343 0.1838
PGSR [8] $\times$ TSDF 0.1670 0.2188 0.1929 0.7622 0.2227
Ours $\times$ TSDF 0.1001 0.1114 0.1057 0.8393 0.4711
@@ 5.1 Settings
For open-vocabulary segmentation, we follow LangSplat [34] to conduct experiments on the LERF-OVS dataset. We adopt the evaluation metrics from Gaussian Grouping. Specifically, we use the text query to select 3D Gaussians, then calculate the average IoU (mIoU) and boundary IoU (mBIoU) accuracy between rendered masks and annotated object masks. For indoor surface reconstruction, we conduct experiments on two RGBD datasets: MuSHRoom [36] and ScanNet++ [49]. We adopt PSNR, SSIM and LPIPS metrics for rendered images and follow DN-Splatter [40] to evaluate mesh quality through five metrics: Accuracy, Completion, Chamfer-L1 distance, Normal Consistency and F-scores. For novel view synthesis, we follow standard PSNR, SSIM and LPIPS metrics for rendered images.
@@
We compare our model against a series of baseline approaches on two tasks. a) open-vocabulary segmentation: Langsplat [34], Gaussian Grouping [47] and OpenGaussian [43]. We deploy the scheme of OpenGaussian to render the selected objects. Specifically, we compute the similarity between 3D Gaussian semantic features and text features, then select Gaussians with high similarity and obtain the rendered object by the rasterizer. b) Indoor surface reconstruction: MonoSDF [51], 2DGS [20], PGSR [8] and DN-Splatter [40]. To align with the setting of DN-Splatter, we introduce the sensor depth to supervise $D_{p}$ by Mean Absolute Error loss.
@@
Our code is built based on PGSR [8]. The densification strategy is adopted from AbsGS [48]. We train GLS for 30k iterations, consuming about 40 minutes on a single NVIDIA RTX 4090 GPU. For indoor surface reconstruction, we first generate rendered depth in each training view, followed by performing the TSDF Fusion [32, 29] to extract the mesh in the TSDF field. Subsequently, for open-vocabulary segmentation, we first compute the similarity between 3D Gaussian semantic features and text features, then filter Gaussians with high similarity to render and reconstruct the selected object.

### v2

@@ Abstract
Recently, 3D Gaussian Splatting (3DGS) has achieved impressive performance on indoor surface reconstruction and 3D open-vocabulary segmentation. This paper presents GLS, a unified framework of 3D surface reconstruction and open-vocabulary segmentation based on 3DGS. GLS extends two fields by improving their sharpness and smoothness. For indoor surface reconstruction, we introduce surface normal prior as a geometric cue to guide the rendered normal, and use the normal error to optimize the rendered depth. For 3D open-vocabulary segmentation, we employ 2D CLIP features to guide instance features and enhance the surface smoothness, then utilize DEVA masks to maintain their view consistency. Extensive experiments demonstrate the effectiveness of jointly optimizing surface reconstruction and 3D open-vocabulary segmentation, where GLS surpasses state-of-the-art approaches of each task on MuSHRoom, ScanNet++ and LERF-OVS datasets. Project webpage: https://jiaxiongq.github.io/GLS_ProjectPage.
@@ 1 Introduction
Surface reconstruction[22, 24, 58, 57, 10, 59] and 3D open-vocabulary segmentation[39, 44, 60, 52, 23, 49] based on 3DGS[29], being widely applied in AR/VR[27] and embodied intelligence[8, 38] due to its capabilities of efficient training and real-time rendering. Recently, notable works have achieved significant progress in both areas. For surface reconstruction, SuGaR[22] proposes regularization terms to align Gaussians and scene surface, then uses Possion reconstruction[28] to extract mesh from Gaussians. For 3D open-vocabulary segmentation, LangSplat[39] and OpenGaussian[49] successfully introduces SAM[31] and CLIP[26] to 3DGS. Gaussian grouping[52] utilizes a universal temporal propagation model DEVA[14] to obtain consistent object masks across views and propose a 3D regularization term to grouping Gaussians. However, these methods only focus on one task and suffer from unstable performance in complex indoor scenes as Fig. 1 shows.
@@
Figure 1: Indoor surface reconstruction and 3D open-vocabulary segmentation. Shadow and high-light regions make state-of-the-art methods struggle in indoor scenes. Our proposed GLS jointly optimizes two tasks based on 3DGS and achieves much better results than OpenGaussian[49] and 2DGS[24].
@@
In this work, we aim to achieve efficient and robust indoor surface reconstruction and 3D open-vocabulary segmentation based on 3DGS. Our goal has two main motivations. On the one hand, the 2D open-vocabulary supervision[31, 26] is naturally view-inconsistent, which easily results in noise body and blur boundary of the segmented object from Gaussians. The accurate scene surface consists of sharp and smooth object surfaces, illustrating that Gaussians are mainly distributed on objects and preserve the object boundaries. This property can make the object segmentation results from Gaussians cleaner and sharper. On the other hand, due to the complex materials and lighting conditions of the indoor scene, the shadow and high-light regions on texture-less and reflective objects always result in noisy surfaces. Fortunately, accurate object masks erase the interference details on the object’s surface. Hence, the object segmentation results can supply the smoothness prior to reducing the reconstruction noises of these objects. In general, the optimization goals of the two tasks can be considered to be the same.
@@
Based on above motivations, we introduce GLS, which leverages complementary between surface reconstruction and 3D open-vocabulary segmentation to boost the performance of 3DGS in two tasks. The framework is presented in Fig. 3. Specifically, we first introduce the normal prior[3] to regularize the surface normal estimated from rendered depth. Then we analyzed different situations of rendered depth under different normal errors and propose a regularization term to enhance the sharpness of the scene surface. To integrate the open-vocabulary information, we add Gaussian semantic features into vanilla 3DGS and then utilize the consistent object masks from DEVA and image features from CLIP to supervise them. In addition, we consider the clip features as the smoothness prior to strengthen the accuracy of texture-less and reflective surfaces. Finally, we adopt TSDF fusion[37] from rendered depth to extract scene mesh, then compute the similarity between input text embeddings and learned semantic embeddings to acquire object masks. We conduct extensive experiments on MuSHRoom[41], ScanNet++[54] and LERF-OVS [39, 30] datasets. As Fig. 2 shows, the superior performance of our model on both tasks demonstrates the effectiveness of connecting indoor surface reconstruction and 3D open-vocabulary segmentation in 3DGS. GLS makes the reconstructed scene surface interactive. The application demos can be visualized in the supplementary videos. In summary, our technical contributions can be listed as follows:
@@
    We design a novel framework based on 3DGS, by jointly optimizing surface reconstruction and 3D open-vocabulary segmentation in complex indoor scenes.
@@
    We propose two novel regularization terms with the help of geometric and semantic cues, to facilitate the sharpness and smoothness of reconstructed scene surfaces and segmented objects.
@@
    Our method inherits the training and rendering efficiency of 3DGS and achieves state-of-the-art accuracy on surface reconstruction and 3D open-vocabulary segmentation tasks.
Figure 2: Effect of joint optimization between indoor surface reconstruction and 3D open-vocabulary segmentation. Our method significantly improves the quality of Gaussian semantic features and surface reconstruction.
Figure 3: Overview of our framework. We adopt geometric and semantic cues produced by generalizable models to jointly strengthen the reconstruction and segmentation quality of 3DGS. Semantic cues consist of view-consistent segmentation results of DEVA[14] and object-level CLIP features generated from LangSplat[39]. Geometric cue comes from DSINE[3]. Our framework achieves accurate indoor surface reconstruction and 3D open-vocabulary segmentation via effective regularization terms, which faithfully enable Gaussian primitives to distribute along the object surface.
@@ 2.1 Neural Rendering
Neural Radiance Fields (NeRF)[35] based on implicit representations have achieved remarkable advancements in novel view synthesis. Some NeRF-based approaches[4, 5, 6] concentrate on extending NeRF to more challenging scenes, while they are still limited by low training and rendering efficiency. Alternatively, other NeRF-based methods[45, 17, 40, 36, 55, 13] utilize explicit representations such as voxels, grids, point clouds and meshes. They faithfully reduce the large computational cost of neural networks. The recent technologies[29, 56, 33] of neural rendering based on alpha-blend rendering have further advanced the rendering speed and quality, by optimizing the attributes of the explicit 3D Gaussians. Our framework is built upon 3DGS and focuses on optimizing it to tackle both surface reconstruction and 3D open-vocabulary segmentation tasks.
@@
Surface reconstruction from multi-view images[43, 19, 20, 21, 42, 57] is a challenging task in computer vision and graphics. Recently, extracting scene surfaces from Gaussian primitives has become a popular research topic. SuGaR[22] utilizes Poisson reconstruction[28] to extract mesh from sampled Gaussians and encourages Gaussians attach to the scene surface by several regularization terms, for enhancing reconstruction accuracy. However, due to the disorder of the Gaussians and the complexity of the scene, fitting Gaussians on geometric surfaces is challenging. 2DGS[24] replaces 3D Gaussian primitives with 2D Gaussian disks for efficient training. They also introduce two regularization terms of depth distortion and normal consistency to reconstruct a smooth scene surface. PGSR[10] estimates the unbiased depth to disable the degeneration of the blending coefficient, and introduces the multi-view regularization term to achieve global-consistent reconstruction. DN-Splatter[46] and VCR-GauS[11] leverage normal or depth priors from generalizeable models[3, 7, 18], to learn effective surfel representations. FDS[12] introduce the optical flow into 3DGS for more accurate geometry.
Due to the lack of object attributes, these methods struggle to reconstruct sharp objects against complex lighting conditions of an indoor scene. To address these issues, we propose combining 3D open-vocabulary segmentation and surface reconstruction. In contrast with previous related works[51, 47, 48], our method releases the need for ground-truth segmentation supervision and achieves highly efficient training speed.
@@ 4.1 Leveraging 2D Semantic and Geometric Cues
For 3D open-vocabulary segmentation, we directly use the view-consistent segmentation results of DEVA[14] as the supervision of the rendered mask $M$. Specifically, given the rendered features $F$, we use an MLP layer to increase its feature dimension to the number of total categories first. The softmax function and standard cross-entropy loss $\mathcal{L}_{m}$ are adopted for final classification, which is defined as: $\mathcal{L}_{m}=-\sum_{i=1}^{S}y_{i}\log(\hat{y}_{i})$, where $S$ is the number of classes, $y_{i}$ is the true label (0 or 1) for the $i$-th class, and $\hat{y}^{i}$ is the predicted probability for the $i$-th class. To enhance the open-vocabulary capabilities of our model, we employ the CLIP features $\hat{F}$ to supervise $F$ by:
@@ 4.2 Semantic-feature Guided Normal Smoothing
Big surfaces like floors and desktops are generally over-smoothing in an indoor scene. Although $\mathcal{L}_{n}$ supply local smoothness for reconstruction, it still struggles in shadow and high-light regions of these surfaces. High-weight $\mathcal{L}_{n}$ causes an over-smoothing effect and then makes some small objects disappear, while low-weight $\mathcal{L}_{n}$ disables the local smoothness of rendered depth.
To resolve this issue, we propose introducing clip features to help smooth big surfaces, then we can employ $\mathcal{L}_{n}$ with low weight for protecting small object boundaries. Essentially, clip features supply high-dimension object-level smoothness which can implicitly reduce the inference of big surfaces. Concretely, we first introduce SAM to select objects that occupy the top-$k$ area by the mask $M_{o}$. and design a novel regularization term $\mathcal{L}_{s}$ for smoothing them as:
@@ 4.4 Optimization
We set $\alpha_{n}=0.07$, $\alpha_{m}=0.3$, $\alpha_{clip}=1.0$, $\alpha_{d}=0.01$ and $\alpha_{s}=0.5$ in our experiments.
Table 1: Quantitative results of indoor surface reconstruction on MuSHRoom dataset (only contains: ‘coffee_room’, ‘computer’, ‘honka’, ‘kokko’ and ‘vr_room’). Our method outperforms other methods on most metrics. The best metrics are highlighted.
@@
DN-Splatter[46] 0.0402 0.0211 0.0306 0.8449 0.8590 1.0h
Ours 0.0288 0.0254 0.0271 0.8640 0.8924 0.7h
2DGS[24] $\times$ 0.1728 0.1619 0.1674 0.7113 0.3433 0.3h
PGSR[10] $\times$ 0.4957 0.7672 0.6315 0.5566 0.1178 0.7h
FDS[12] $\times$ 0.0639 0.0528 0.0584 0.8169 0.6998 1.3h
Ours $\times$ 0.0538 0.0582 0.0560 0.8357 0.6922 0.7h
Table 2: Quantitative results of indoor surface reconstruction on ScanNet++ dataset (only contains: ‘8b5caf3398’ and ‘b20a261fdf’). The best metrics are highlighted.
@@
DN-Splatter[46] 0.0977 0.0431 0.0704 0.8272 0.7094 1.0h
Ours 0.0640 0.0272 0.0444 0.9064 0.8623 0.6h
2DGS[24] $\times$ 0.2440 0.4362 0.3401 0.6343 0.1838 0.2h
PGSR[10] $\times$ 0.1670 0.2188 0.1929 0.7622 0.2227 0.6h
Ours $\times$ 0.0861 0.1009 0.0935 0.8578 0.4799 0.6h
@@ 5.1 Settings
For 3D open-vocabulary segmentation, we follow LangSplat[39] to conduct experiments on the LERF-OVS dataset[30]. We adopt the evaluation metrics from Gaussian Grouping. Specifically, we use the text query to select 3D Gaussians, then calculate the average IoU (mIoU) and boundary IoU (mBIoU) accuracy between rendered masks and annotated object masks. For indoor surface reconstruction, we conduct experiments on two RGBD datasets: MuSHRoom[41] (only contains: ‘coffee_room’, ‘computer’, ‘honka’, ‘kokko’ and ‘vr_room’) and ScanNet++[54] (only contains: ‘8b5caf3398’ and ‘b20a261fdf’). We and the same tool of DN-Splatter[46] to evaluate mesh quality through five metrics: Accuracy, Completion, Chamfer-L1 distance, Normal Consistency and F-scores. For novel view synthesis, we follow standard PSNR, SSIM and LPIPS metrics for rendered images.
@@
We compare our model against a series of baseline approaches on two tasks. a) 3D open-vocabulary segmentation: Langsplat[39], Gaussian Grouping[52] and OpenGaussian[49]. We deploy the scheme of OpenGaussian to render the selected objects. Specifically, we compute the similarity between 3D Gaussian semantic features and text features, then select Gaussians with high similarity and obtain the rendered object by the rasterizer. b) 3DGS-based indoor surface reconstruction: 2DGS[24], PGSR[10], DN-Splatter[46] and FDS[12]. For fair comparison among all methods, we take the LiDAR points of iPhone as the default input point cloud on both datasets, and re-ran their source codes released in GitHub in 5 scenes. To align with the setting of DN-Splatter, we introduce the sensor depth to supervise $D_{p}$ by Mean Absolute Error loss.
@@
Our code is built based on PGSR[10]. The densification strategy is adopted from AbsGS[53]. We train GLS for 30k iterations, consuming about 40 minutes on a single NVIDIA RTX 4090 GPU. For indoor surface reconstruction, we first generate rendered depth in each training view, followed by performing the TSDF Fusion[37, 34] to extract the mesh in the TSDF field. Subsequently, for 3D open-vocabulary segmentation, we first compute the similarity between 3D Gaussian semantic features and text features, then filter Gaussians with high similarity to render and reconstruct the selected object. More details can be seen in the supplementary materials.